Write the rate of as a fraction in lowest term
32 cups for 6 people's
The rate is people's/cups
Or the rate it cups over people

Answer by addingup(3677)

you have more cups per people which means everyone gets more than one cup. So you have to divide your cups over the number of people like this 32/6 = 5.33 you can't break a cup, so 4 people will get 5 cups and 2 people will get 6 cups:
4*5= 20; 20+(2*6) = 32 so you have now distributed your 32 cups over 6 people